(a) worked out incorrectly; or (b) worked out correctly but using a method not specified in the problem statement; or (c) not2) Write the nth term of the sequence whose first several terms are -12, -9, -6, -3, ... (4 points)
Answer #1

Given, sequence, 18:- -12,-9.-6, -3, -- solution : Given First term, a=-12 d=-12+127-6-64)= -3-461-3 since, Common difference